OpenStudy (anonymous):

What is 3 to the power of 3 over 2 equal to? cube root of 9 square root of 9 cube root of 27 square root of 27 @iGreen.

OpenStudy (anonymous):

3^(3/2) = 5.19615242 See which other option equals 5.19615242..

OpenStudy (anonymous):

none

OpenStudy (anonymous):

its not showing the correct anser choices up there

OpenStudy (anonymous):



OpenStudy (anonymous):

Look at this rule: \(x^{\frac{m}{n}} =~^n \! \! \! \! \! \sqrt{x^m}\)

OpenStudy (anonymous):